# Collective Outlier Detection and Enumeration with Conformalized Closed Testing

### v2

@@ 13      $d_{\mathrm{Shiraishi}}(S)=d_{\mathrm{Shiraishi}}(S)+1$;
A4.1 Numerical Experiments with Synthetic Data
Figures A1 and A2 describe findings similar to those in Figures 2 and 3, respectively. The distinction is that now the data are simulated from a binomial model borrowed from Liang et al. (2024), for which binary classifiers are more powerful than one-class classifiers. Figures A3 and A4 summarize the 90-th empirical quantiles of the lower confidence bounds for the numbers of outliers presented in Figures 2 and A1, respectively. Figure A5 delves into experiments related to those depicted in Figures A1 and A2, but utilizing synthetic data from a mixture model that bridges between the distributions considered above. These results highlight ACODE’s flexibility, which selects an effective classifier and local testing procedure for each specific case.
Figures A6 and A7 offer additional insights into the performance of the LMPI local testing procedure, outlined in Section 2.3.2, when applied in different settings. These results highlight its practical advantages in situations where the outlier scores are either underdispersed or overdispersed relative to the inlier scores.
@@
Figure A2: Lower confidence bounds for the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. In these experiments, ACODE is applied using a one-class support vector classifier to compute the conformity scores.
@@
Figure A6: Power of different rank tests for the global null hypothesis of no outliers in a test set containing 200 independent observations of a simulated univariate score, using a reference set containing 500 independent inlier scores. The nominal level is $\alpha=0.05$ (horizontal dashed line). The results are shown as a function of the true number of outliers in the test set. Top: the inliers are randomly sampled from a uniform distribution on $[0,1]$, while the outliers are sampled from the alternative distribution indicated in each panel. Bottom: the inliers are standard normal, while the outlier scores follow a non-standard normal distribution, as indicated in each panel. Three versions of the LMPI test from Section 2.3.2 are compared. The first version uses oracle knowledge of the true transformation $G$ linking the outlier distribution to the inlier distribution. The second version of the LMPI test uses an empirical estimate of $G$ obtained as described in Appendix A2. The third version relies on a monotone approximation of the derivative $g=G^{\prime}$, which (in Figure A7) enables a computationally convenient shortcut in the context of closed testing, as also explained in Appendix A2.
Figure A7: Median values for a 90% lower confidence bound on the total number of outliers in the test set, in the same experiments of Figure A6. The lower bound is calculated through closed testing, using different local testing procedures. For the LMPI approach, if the derivative $g$ of the oracle function $G$ is not monotone, it is in practice replaced by a monotone approximation that enables the application of a computationally efficient closed testing shortcut.
A4.2 Numerical Experiments with Particle Collision Data
Figure A8 is a more detailed version of Figure 1, including a broader range for the choice of the local test. Figures A9 and A10 report on experiments similar to those of Figures 1 and 5, respectively, but leveraging a larger training set containing 100,000 inliers, which leads to higher power for all methods. Tables A3 and A4 provide a more detailed view of the results, respectively, in Figures A8 and A9. Finally, Figures A11 and A12 summarize the 90-th empirical quantiles of the lower confidence bounds for the numbers of outliers presented in Figures 5 and A10, respectively, confirming again the validity of our method.
@@
Outliers Simes Storey-Simes Fisher WMW LMPI (Lehmann, k=3) Adaptive Cherry picking
@@
Table A3: Performance of ACODE for collective outlier detection with the LHCO data, as a function of the true number of outliers. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). The numbers in parenthesis are standard errors. Other details are as in Figures 1 and A8.
Figure A9: Performance of ACODE for collective outlier detection with the LHCO data, using a training set containing 100,000 inliers. In the interest of computational efficiency, in these experiments, the conformity scores leveraged by ACODE are computed by a fixed AdaBoost classifier. Other details are as in Figure 1.
Local testing procedure
Outliers Simes Storey-Simes Fisher WMW LMPI (Lehmann, k=3) Adaptive Cherry picking
90% Lower bound (median)
   0 0 (0) 0 (0) 0 (1) 0 (3) 0 (2) 0 (2) 0 (3)
   200 0 (0) 0 (0) 0 (4) 0 (7) 0 (7) 0 (7) 0 (8)
   500 0 (0) 0 (0) 204 (11) 202 (16) 247 (15) 273 (15) 284 (15)
   1000 88 (10) 118 (10) 666 (10) 720 (18) 751 (14) 758 (13) 796 (13)
   1500 343 (16) 392 (15) 1114 (10) 1224 (20) 1242 (16) 1252 (15) 1276 (15)
90% Lower bound (90-th quantile)
   0 0 (0) 0 (0) 0 (1) 0 (3) 0 (2) 2 (2) 8 (3)
   200 0 (0) 0 (0) 46 (4) 107 (7) 131 (7) 133 (7) 156 (8)
   500 0 (0) 0 (0) 328 (11) 421 (16) 436 (15) 456 (15) 474 (15)
   1000 238 (10) 251 (10) 782 (10) 952 (18) 919 (14) 932 (13) 957 (13)
   1500 541 (16) 563 (15) 1255 (10) 1467 (20) 1430 (16) 1452 (15) 1478 (15)
Power (global null)
   0 0.00 (0.00) 0.00 (0.00) 0.04 (0.02) 0.06 (0.02) 0.07 (0.03) 0.11 (0.03) 0.14 (0.03)
   200 0.00 (0.00) 0.00 (0.00) 0.28 (0.05) 0.23 (0.04) 0.31 (0.05) 0.35 (0.05) 0.40 (0.05)
   500 0.06 (0.02) 0.07 (0.03) 0.89 (0.03) 0.78 (0.04) 0.84 (0.04) 0.88 (0.03) 0.89 (0.03)
   1000 0.76 (0.04) 0.83 (0.04) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
   1500 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Table A4: Performance of ACODE for collective outlier detection with the LHCO data, using a training set containing 100,000 inliers. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis (bottom). The numbers in parenthesis are standard errors. Other details are as in Figure A9.
Figure A10: Lower confidence bounds for the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. The dashed curve corresponds to the true number of outliers in this selected set. In these experiments, ACODE is applied using a one-class support vector classifier to compute the conformity scores.
Figure A11: Empirical 90-th quantile for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in the same numerical experiments of Figure 5.
Figure A12: Empirical 90-th quantile for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in the same numerical experiments of Figure A10.
A4.3 Additional Experiments with Real Data
In this section, we further investigate the performance of ACODE by applying it to 6 additional data sets previously utilized in the related literature. These data sets are:
  • CreditCard (Dataset, c), containing 30 variables, 492 outliers, 284315 inliers;
  • Pendigits (Dataset, e), containing 16 variables, 156 outliers, 6714 inliers;
  • Covertype (Dataset, b), containing 10 variables, 2747 outliers, 286048 inliers;
  • Shuttle (Dataset, f), containing 9 variables, 3511 outliers, 45586 inliers;
  • Mammography (Dataset, d), containing 6 variables, 260 outliers, 10923 inliers;
  • ALOI (Dataset, a), containing 27 variables, 1508 outliers, 48026 inliers.
@@
Figures A13A18 summarize, separately for each data set, the performances of the lower confidence bounds and global tests obtained with our method. The results are reported as a function of the true number of outliers in the test set and are also stratified based on the type of classifier utilized by ACODE and on the underlying local testing procedure. The findings overall indicate that binary classification models generally yield higher power in these experiments, which can be attributed to the relatively low dimensionality of these data sets. Moreover, among the local testing methods considered, both Fisher’s method and the WMW method demonstrate similarly high power, typically outperforming Simes’ method. Crucially, the fully automatic implementation of ACODE, which selects both the classification model and the local testing procedure in a data-adaptive way, leads to near-oracle performance in all cases.
Finally, Figures A19 and A20 describe experiments aimed at constructing 90% lower confidence bounds for the number of outliers within a data-dependent subset of test points, specifically those selected for their high conformity scores. For easier understanding of these experiments, ACODE is now applied with a fixed classification algorithm, the one-class isolation forest, rather than leveraging the full suite of 6 different algorithms. The results show that the WMW test consistently outperforms the Fisher combination method in terms of power, aligning with findings previously documented in Sections 3.1 and 3.2. Furthermore, the fully adaptive implementation of ACODE effectively maximizes power by automatically identifying the most powerful local testing procedure for each case, again achieving oracle-like performance.

### v3

@@ Results.
A6.1 Numerical Experiments with Synthetic Data
Figure A2 illustrates the effect of the proportion of calibration data allocated to tuning on ACODE’s performance. Specifically, we repeat the experiments summarized in Figure 2 while varying the fraction of calibration points used for tuning. The total calibration sample size is fixed at 1000, of which 250 points (25%) were used for tuning in Figure 2; here we vary this proportion from 0.01 to 0.99. The results show that ACODE’s performance remains quite stable over a wide range of tuning fractions, but degrades when too few points are used for tuning (rendering tuning ineffective) or when too many are allocated to tuning (leaving insufficient data for computing significance tests). This motivates our choice of 25% as a reasonable default.
Figure A2: Median values for a 90% lower confidence bound on the number of outliers in a test set, computed by ACODE on synthetic data based on different classifiers and local testing procedures. The results are shown as a function of the proportion of calibration points (1000 observations in total) used for tuning (ref. Section 2.6), for different true numbers of outliers within a test set of size 1000. Other details are as in Figure 2, where the proportion of calibration data points used for tuning is fixed at 0.25.
Figure A3 provides additional details on the experiments of Figure 3, reporting the empirical 90-th quantiles for the 90% lower confidence bounds, instead of the median.
Figure A3: Empirical 90-th quantile for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in experiments similar to those of Figure 2. Other details are as in Figure 3.
To assess the difficulty of individual outlier detection in this synthetic setting, we additionally compare ACODE with the Benjamini-Hochberg (BH) (Benjamini and Hochberg, 1995) procedure applied to individual conformal $p$-values at nominal FDR level 10%, following the approach of Bates et al. (2023). As shown in Figure A4, individual discovery is feasible in this setting, due to the relatively strong signals, but consistently less powerful than collective outlier enumeration with ACODE.
Figure A4: Median values for a 90% lower confidence bound on the number of outliers in a test set, computed by ACODE on synthetic data based on different classifiers and local testing procedures, as in Figure 2. Here, we introduce an additional benchmark: the number of discoveries made by BH procedure for individual outlier identification with FDR control, applied at nominal FDR level 10%. This shows our ACODE method has higher power, except when applied using Simes’ test as local testing procedure.
Figures A5 and A6 describe findings similar to those in Figures 2 and 3, respectively. The distinction is that now the data are simulated from a binomial model borrowed from Liang et al. (2024), for which binary classifiers are more powerful than one-class classifiers. Figures A7, A8, and A9 summarize the 90-th empirical quantiles of the lower confidence bounds for the numbers of outliers presented in Figures 2, A5, and A6, respectively. Figure A10 delves into experiments related to those depicted in Figures A5 and A6, but utilizing synthetic data from a mixture model that bridges between the distributions considered above. These results highlight ACODE’s flexibility, which selects an effective classifier and local testing procedure for each specific case.
Figures A11 and A12 offer additional insights into the performance of the Shiraishi local testing procedure, outlined in Section 2.3.2, when applied in different settings. These results highlight its practical advantages in situations where the outlier scores are either underdispersed or overdispersed relative to the inlier scores.
@@
Figure A6: Median lower confidence bounds for the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers in the test set. The dashed curve indicates the true number of outliers in the selected subset, averaged over 100 independent experiments (the true count varies across runs, though with relatively low variance). In these experiments, ACODE is applied using a one-class support vector classifier to compute the conformity scores.
@@
Figure A8: Empirical 90-th quantile for a 90% lower confidence bound for the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments otherwise similar to those of Figure 2. Other details are as in Figure A6.
@@
Figure A11: Power of different rank tests for the global null hypothesis of no outliers in a test set containing 200 independent observations of a simulated univariate score, using a calibration set containing 500 independent inlier scores. The nominal level is $\alpha=0.05$ (horizontal dashed line). The results are shown as a function of the true number of outliers in the test set. Top: the inliers are randomly sampled from a uniform distribution on $[0,1]$, while the outliers are sampled from the alternative distribution indicated in each panel. Bottom: the inliers are standard normal, while the outlier scores follow a non-standard normal distribution, as indicated in each panel. Three versions of the Shiraishi test from Section 2.3.2 are compared. The first version uses oracle knowledge of the true transformation $G$ linking the outlier distribution to the inlier distribution. The second version of the Shiraishi test uses an empirical estimate of $G$ obtained as described in Appendix A2. The third version relies on a monotone approximation of the derivative $g=G^{\prime}$, which (in Figure A12) enables a computationally convenient shortcut in the context of closed testing, as also explained in Appendix A2.
Figure A12: Median values for a 90% lower confidence bound on the total number of outliers in the test set, in the same experiments of Figure A11. The lower bound is calculated through closed testing, using different local testing procedures. For the Shiraishi approach, if the derivative $g$ of the oracle function $G$ is not monotone, it is in practice replaced by a monotone approximation that enables the application of a computationally efficient closed testing shortcut.
A6.2 Numerical Experiments with Particle Collision Data
Figure A13 is a more detailed version of Figure 1, including a broader range for the choice of the local test.
@@
Figure A14 presents results from experiments in which ACODE is used to construct 90% lower confidence bounds for the number of outliers in a data-driven subset of test points, selected as those with the largest conformity scores. To reduce computational cost, we consider smaller test sets of 2000 data points with varying proportions of outliers, and we do not apply ACODE with the Shiraishi local test, whose implementation becomes relatively expensive when the selected set differs from the full test set. As in Figure 3, conformity scores are computed using a single classifier (AdaBoost) to facilitate interpretation. The results confirm that the adaptive version of ACODE again approximately maximizes power by automatically selecting the testing procedure.
Figure A14: Median values over repeated experiments for a 90% lower confidence bound on the number of outliers within an adaptively selected subset of 1000 test points, in numerical experiments similar to those of Figure 2. The results are shown as a function of the proportion of selected test points and of the total number of outliers. The dashed curve indicates the true number of outliers in the selected subset, averaged over 100 independent experiments (the true count varies across runs, though with relatively low variance).
Table A4 provides a more detailed view of the results in Figure A13.
@@
Outliers Simes Storey-Simes Fisher WMW LMPR (Lehmann, k=3) Adaptive Cherry picking
@@
Table A4: Performance of ACODE for collective outlier detection with the LHCO data, as a function of the true number of outliers. The performance is measured in terms of median and 90-th percentile values of a 90% lower confidence bound for the number of outliers (top and center) and the power to reject the global null hypothesis of no outliers (bottom). The numbers in parenthesis are standard errors. Other details are as in Figures 1 and A13.
A6.3 Additional Experiments with Real Data
In this section, we further investigate the performance of ACODE by applying it to 6 additional data sets previously utilized by in the related literature (Bates et al., 2023; Liang et al., 2024; Marandon et al., 2024). These data sets are:
  • ALOI (Dataset, ), containing 27 variables, 1508 outliers, 48026 inliers. Previously used by Bates et al. (2023).
  • Covertype (Dataset, ), containing 10 variables, 2747 outliers, 286048 inliers. Previously used by Bates et al. (2023).
  • CreditCard (Dataset, ), containing 30 variables, 492 outliers, 284315 inliers. Previously used by Bates et al. (2023) and Marandon et al. (2024).
  • Mammography (Dataset, ), containing 6 variables, 260 outliers, 10923 inliers. Previously used by Bates et al. (2023), Liang et al. (2024) and Marandon et al. (2024).
  • Pendigits (Dataset, ), containing 16 variables, 156 outliers, 6714 inliers. Previously used by Bates et al. (2023).
  • Shuttle (Dataset, ), containing 9 variables, 3511 outliers, 45586 inliers. Previously used by Bates et al. (2023) and Marandon et al. (2024).
@@
Figures A15A20 summarize, separately for each data set, the performances of the lower confidence bounds and global tests obtained with our method. The results are reported as a function of the true number of outliers in the test set and are also stratified based on the type of classifier utilized by ACODE and on the underlying local testing procedure. The findings overall indicate that binary classification models generally yield higher power in these experiments, which can be attributed to the relatively low dimensionality of these data sets. Moreover, among the local testing methods considered, both Fisher’s method and the WMW method demonstrate similarly high power, typically outperforming Simes’ method. Crucially, the fully automatic implementation of ACODE, which selects both the classification model and the local testing procedure in a data-adaptive way, leads to near-oracle performance in all cases.
Finally, Figures A21 and A22 describe experiments aimed at constructing 90% lower confidence bounds for the number of outliers within a data-dependent subset of test points, specifically those selected for their high conformity scores. For easier understanding of these experiments, ACODE is now applied with a fixed classification algorithm, the one-class isolation forest, rather than leveraging the full suite of 6 different algorithms. The results show that the WMW test consistently outperforms the Fisher combination method in terms of power, aligning with findings previously documented in Sections 3.1 and 3.2. Furthermore, the fully adaptive implementation of ACODE effectively maximizes power by automatically identifying the most powerful local testing procedure for each case, again achieving oracle-like performance.
@@
A6.4 Ludbrook and Dudley (1998) Example
Analogously to signal detection, where—depending on sparsity and signal strength—a signal may be estimable, detectable, or undetectable (see Fig. 1 of Donoho and Jin (2004)), our problem admits a similar hierarchy. Outliers may be individually identifiable, enumerable (and thus detectable in aggregate), or undetectable. These inferential tasks form a strict ordering: detection is easier than enumeration, which in turn is easier than identification. Less demanding tasks yield higher power but provide coarser information, reflecting an inherent trade-off between power and informativeness.
Our closed testing–based approach navigates this trade-off: when effects are detectable, it enables enumeration; when they cannot be pinpointed, it allows them to be progressively bracketed within subsets of observations.
To illustrate this point, we consider a simple example based on the data of Ludbrook and Dudley (1998). This example shows that even when individual-level detection fails, in particular, when no single $p$-value falls below the significance level $\alpha$, group-level inference can still reveal the presence of outliers. Specifically, the closed testing procedure yields a lower bound on the number of outliers and, , by examining subsets of observations, allows them, even if they cannot be pinpointed individually, to be bracketed within a smaller group of observations.
Ludbrook and Dudley (1998, section 4.3) reports the calibration and test scores:
$(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6},x_{7})$ $=$ $(5.42,5.86,6.16,6.55,6.8,7,7.11),$
$(y_{1},y_{2},y_{3},y_{4},y_{5})$ $=$ $(6.51,7.56,7.61,7.84,11.5).$
The resulting conformal $p$-values are:
$(p_{1},p_{2},p_{3},p_{4},p_{5})=(0.625,0.125,0.125,0.125,0.125).$
At the nominal level $\alpha=0.05$, all $p$-values exceed $\alpha$, and neither the Benjamini–Hochberg nor the Storey procedure rejects any hypotheses. In contrast, the WMW rank test rejects the global null hypothesis that there are no outliers with a $p$-value of 0.03, as reported in Ludbrook and Dudley (1998). This suggests that, although individual outliers may not be identified, a group of outliers may still be detected.
Using the closed testing procedure based on local WMW tests, we obtain an overall bound of $d=2$. This means that, with 95% confidence, at least two of the five test observations are outliers, although we cannot determine exactly which ones.
To narrow down location of the outliers, we can examine subsets of $[n]$. For example, $d(S)=2$ for $S=\{2,3,4,5\}$, indicating that at least two outliers are among the four observations with the largest test scores. Zooming in further, for $S^{\prime}=\{3,4,5\}$ and $S^{\prime\prime}=\{4,5\}$ we obtain $d(S^{\prime})=1$ and $d(S^{\prime\prime})=0$, indicating that while at least one outlier lies among the three highest-scoring test observations, the outliers cannot be bracketed further.
Note that if interest is restricted to an a priori defined subset $R$ of the test observations, then applying the closed testing procedure directly to $R$ is at least as powerful, and typically more powerful, than applying the procedure to the full set $[n]$ and subsequently focusing on $R$. For example, if the set $R=\{2,3,4,5\}$ had been specified a priori as the hypotheses of interest, the resulting lower bound would improve from 2 to 3. Choosing a priori a subset does not always lead to an improvement. For example, the set $\{1,2,3,4\}$ would result in the same lower bound of 1.